log7 + log (n - 2) = log 6n
Answer by stanbon(75887)

log7 + log (n - 2) = log 6n
Log[7(n-2)] = log6n
Take the anti-log to get:
7(n-2) = 6n
7n-14 = 6n
n = 14
==================
Cheers,
Stan H.